OpenStudy (anonymous):

Which values of x and y will satisfy y = 5x and 12x + 9y = 60? A. x = 1, y = 19 B. x = 35 y= 20/19 C. x = 1/20 y = 1/90 D. x = 20/19 y = 100/19 E. x = 19 y = 50/19

OpenStudy (shamil98):

y = 5x 12x + 9y = 60 Use substitution and solve the system 12x + 9(5x) = 60 12x + 45x = 60 57x = 60 x=20/19 y = 5(20/19) y= 100/19